# Does AI Code Review Lead to Code Changes? A Case Study of GitHub Actions

### v1

@@ V Discussion
Does AI-based code review lead to code changes? In general, human-authored review comments are much more likely to be addressed than AI-generated ones: 60% of valid human review comments led to code changes, compared to only 0.9%–19.2% for valid AI-generated comments depending on the tool (Table VIII). However, the answer is not simply “no”. Our results show that there is a clear potential for AI-generated code review to influence development if these tools are designed well.
@@ V-A Design Matters
Through our interpretability analysis (Table X), we identified several characteristics that consistently increased the likelihood of AI-generated comments being addressed.
@@
  • Manual triggering: Tools explicitly triggered by developers produce higher addressing rates. For example, anc95/ChatGPT-CodeReview showed a 12.8% addressing rate for manually triggered comments versus 6.8% for automatically triggered ones (Table XI).
  • Code-rich, concise comments: Comments with a high code-to-text ratio, especially multiline suggestions, are more likely to be acted upon. We observed a clear increase in the addressing rate when more than half of the comment content was code (left in Table XII).
  • Focus on newcomers: Comments were more likely to be addressed when targeting contributors with little prior commit history. For the least experienced developers, the addressing rate was five times higher than for the most experienced (right in Table XII).
Prompt design also plays an important role. For example, anc95/ChatGPT-CodeReview uses a short, generic prompt such as “Below is a code patch, please help me do a brief code review on it. Any bug risks and/or improvement suggestions are welcome:”, which often results in vague comments. In contrast, coderabbitai/ai-pr-reviewer provides a long and detailed prompt (over 400 words), explicitly instructing the model to avoid generalities and provide focused feedback. It also includes logic to suppress comments such as “LGTM!” unless no issues are found, further reducing noise.
@@
It is also worth noting that not all valid comments necessarily need to result in immediate code changes to be useful. Some AI-generated suggestions, even if ultimately not adopted, may still prompt reflection, discussion, or future improvements. For example, a comment raising a non-critical concern might influence a developer’s approach in a subsequent commit or serve as a lightweight form of documentation.
A fundamental issue underlying these problems is the dominant one-in-one-out paradigm used by most tools. That is, each code change automatically leads to a review comment, regardless of whether a comment is warranted. This leads to low-precision output and reviewer fatigue. Instead, we advocate for a multi-stage approach, similar to practices in traditional machine learning pipelines. The tool should first determine whether a change needs review, then decide whether to generate a comment, and finally, where appropriate, offer a code suggestion. This would bring structure and discipline to AI-based review and help ensure that comments are meaningful and justified.
V-B Recommended Features for Effective AI Code Review
@@
  • Provide richer context to the LLM: Supplying the PR title, description, and summaries of changes helps generate more useful review comments.
@@
V-C Implications for AI Code Review Adopters
@@
Clarify the role of AI code review: Even the best-performing AI tool achieved only a 19.2% addressing rate, far below the human reviewers’ 60%. Therefore, for now, AI code review is better framed as a complement to human review, not a replacement. This requires understanding AI code review’s capability boundaries and deploying it in appropriate scenarios. Based on our LDA topic analysis (RQ3), AI reviewers are good at providing defensive programming suggestions, and developers are relatively more responsive to their comments on accessibility enhancements and concurrency control. Accordingly, we recommend positioning AI reviewers as defenders: tightly scope them to this mandate (with targeted prompts and other recommended features), thereby suppressing category-sprawling yet low-quality comments. This allows human reviewers to focus on higher-level issues, such as overall architecture (e.g. “Try sorting at the DB level, it will be 100x faster”), development context (e.g., “Remove this as project requires Python $\geq$ 3.11”), and business logic (e.g., “This will cause issues on ARM64”). Mapping the boundaries of AI code review’s effectiveness is also an ongoing task for researchers.
@@ VI Threats to Validity
Internal Validity: In RQ2, we extracted 150 comments from over 22,000 AI-generated and 1,166 human-authored comments and obtained gold-standard labels via multi-rater annotation. To assess representativeness, the first author examined an additional 250 samples and observed similar addressing patterns as reported in our main findings. Therefore, we believe the gold standard reflects the broader dataset. For RQ3, our addressing analysis focuses on four popular actions, which may limit the generalizability of our findings. However, within the scope of our dataset, usage was highly concentrated on these four actions, making them representative of real-world deployments observed in our study. Moreover, the final analysis uses a binary classification of whether comments were addressed, which does not capture the degree or impact of the resulting code changes. Nevertheless, our evaluation shows substantial inter-rater agreement and strong LLM performance on this task, ensuring the reliability of our findings. Our feature engineering for impact factor analysis may also be incomplete or miss latent feature interactions. However, our features was comprehensively designed across four dimensions and inspired by prior works [9, 23]. Another threat is that different interpretation methods may yield different analysis results. Since our feature set spans four dimensions, we chose Random Forest to better model the complex and non-linear interactions between features, rather than the simpler model like logistic regression, which assumes a linear relationship between features. We also tried logistic regression in our experiments and it achieved a lower performance (84.8% accuracy and 0.819 Macro-F1) than our Random Forest model (88.5% and 0.854 respectively). This model choice also aligns with work [9, 23]. And the use of SHAP is a standard approach for interpreting such models as well. Therefore, our interpretation pipeline is well-justified. We also acknowledge that these interpretations describe associations between features and the predicted probability of ”Addressed”, not causal effects. If causal identification is desired, further studies (e.g., controlled experiments or developer interviews) would be needed.
External Validity: Our study aims to investigate how comment-related factors influence whether review comments lead to subsequent code changes. To mitigate the impact of inactive repositories which may not yet have established stable review workflows, we followed prior research [19] and collected data only from repositories with at least 50 PRs. Nevertheless, this criterion does not fully capture project scale. We find that most comments in our addressing dataset originate from small- to medium-sized projects (primarily those with $\leq$ 50 non-bot contributors). Therefore, our findings may not generalize to very large-scale projects. Furthermore, our analysis focused on English-language comments. As described in RQ2 Phase I, we began with 16,762 comments from 100 repositories. Applying a language filter reduced this to 4,229 comments from 53 repositories, resulting in an exclusion of approximately 75% of comments and 47 repositories. Among them, 32 repositories were excluded because they only contained Korean comments (e.g., 1,828 comments from KEA-DoKebi/dalkom-backend). Nevertheless, we still retained data from over half of the original projects. And given that English serves as the main communication language in the global development activities [28], with even non-native speakers commonly adopting it, we believe our findings remain relevant to the broader open-source community.
Conclusion Validity: We completed our data collection in early February 2025. Given the rapid evolution of AI-based review tools, our findings may not fully reflect the current landscape, as future updates or tools could differ. To mitigate this threat, we performed a new search on July 23, 2025, and found no significant changes in the top-ranked actions. The only notable newcomer, hustcer/deepseek-review (ranked #7, 324 stars), had its first commit on January 29, 2025, after our action selection on January 7, 2025. We are therefore confident that our sample of GitHub Actions remains representative of the current real-world usage for code review.
@@ VII Related Work
Although some studies include automatic and human evaluations of generated comments [42], most of this research has focused on comment quality rather than downstream impact. Our study complements these efforts by analyzing how AI-generated comments are used in real development workflows and whether they lead to changes.
@@ VIII Conclusion and Future Work
This paper presents a large-scale empirical analysis of AI-based code review actions on GitHub. We examine how these tools are adopted, whether their comments lead to code changes, and what factors influence their effectiveness. Our findings reveal that adoption is highly concentrated on four popular actions. Many AI-generated comments are not addressed, especially when they are vague or lack context. However, comments that are concise, specific are more likely to be addressed—particularly by less experienced contributors. Based on our findings, we offer implications for code review tool designers, adopters, and GitHub.
We recommend treating AI code reviewers as collaborators with human reviewers. To enhance its utility, future work should focus on improving review quality through exploring richer contextual information, designing better prompts, and adopting multi-stage review strategies to generate more specific and actionable feedback. It is also essential to further clarify its capabilities and boundaries to identify optimal ways for AI and human reviewers to collaborate effectively. We hope our findings contribute to the development of more practical and developer-friendly AI-based code review tools.

### v2

@@ V Discussion
Does AI-based code review lead to code changes? In general, human-authored review comments are much more likely to be addressed than AI-generated ones: 60% of valid human review comments led to code changes, compared to only 0.9%–19.2% for valid AI-generated comments depending on the tool (Table IX). However, the answer is not simply “no”. Our results show that there is a clear potential for AI-generated code review to influence development if these tools are designed well.
@@ V-A Design Matters
Through our interpretability analysis (Table XI), we identified several characteristics that consistently increased the likelihood of AI-generated comments being addressed.
@@
  • Manual triggering: Tools explicitly triggered by developers produce higher addressing rates. For example, anc95/ChatGPT-CodeReview showed a 12.8% addressing rate for manually triggered comments versus 6.8% for automatically triggered ones (Table XVI).
  • Code-rich, concise comments: Comments with a high code-to-text ratio, especially multiline suggestions, are more likely to be acted upon. We observed a clear increase in the addressing rate when more than half of the comment content was code (Table XVI).
  • Focus on newcomers: Comments were more likely to be addressed when targeting contributors with little prior commit history. For the least experienced developers, the addressing rate was five times higher than for the most experienced (Table XVI).
To further investigate the effectiveness of review granularity, we randomly investigated 14.4%, i.e., 30 comments, of the 208 PR-level review comments as follows: 5 from Action ID-8, 5 from ID-11, and 20 from ID-16. Only 1 comment was partially addressed, which provided a specific performance improvement: “For better performance, preallocate the ‘architectures’ slice by using ‘make([]types.Architecture, 0, len(fm.Spec.Architectures))’ ”. This aligns with our finding that concrete code suggestions are more likely to be adopted. However, the majority of PR-level comments were ignored, supporting our conclusion that fine-grained (i.e., hunk-level) reviews are more effective for AI-generated code review.
We also highlight that prompt design plays an important role in generating high-quality comments. For example, anc95/ChatGPT-CodeReview uses a short, generic prompt such as “Below is a code patch, please help me do a brief code review on it. Any bug risks and/or improvement suggestions are welcome:”, which often results in vague comments. In contrast, coderabbitai/ai-pr-reviewer provides a long and detailed prompt (over 400 words), explicitly instructing the model to avoid generalities and provide focused feedback. It also includes logic to suppress comments such as “LGTM!” unless no issues are found, further reducing noise.
@@
Figure 4: Illustrative case (link): Effective tool design (hunk-level review, concrete code suggestions, and interactive dialogue) successfully persuading a senior developer to adopt a change.
@@
We present an illustrative case in Figure 4 to show why these design choices matter. This case drawn from coderabbitai/ai-pr-reviewer, demonstrates how concrete modifications and interactive feedback can persuade even an experienced developer (Author_Prior_Commits=186) to accept a suggestion they initially rejected. After the developer identified potential issues with rounding errors, the AI adapted its proposal from a strict equality check to a robust fuzzy comparison, successfully transforming the initial rejection into an actionable improvement.
Beyond post-comment refinement, we also argue that AI tools could be more cautious before posting. We found that most current tools rely on a simplistic one-in-one-out paradigm, where each code change automatically leads to a review comment, regardless of whether a comment is warranted. This leads to low-precision output and reviewer fatigue. Instead, we advocate for a multi-stage approach, similar to practices in traditional machine learning pipelines. The tool should first determine whether a change needs review, then decide whether to generate a comment, and finally, where appropriate, offer a code suggestion. This would bring structure and discipline to AI-based review and help ensure that comments are meaningful and justified.
V-B Beyond Code Changes: Impact on Closed PRs and Feedback on Unaddressed Comments
While we mainly focused on merged PRs to observe definitive code improvements, we recognize that AI-generated review comments can provide value even when they do not lead to a merge or any direct code changes. Therefore, we conducted two supplementary analyses on the comments outside our main dataset.
V-B1 Analysis of AI Comment Impact in Closed PRs
We extended our analysis to closed (non-merged) pull requests for the four primary actions to investigate if developers engage with AI feedback before a PR is abandoned. Following the methodology proposed in RQ2, we applied the LLM-assisted framework to assess whether comments were valid and addressed in the latest commits of these closed PRs. We observed a significantly lower addressing rate for AI comments in closed PRs compared to merged ones: only 3.0% (17/570) for Action ID-1, and 0% for Action ID-2 (n=14n\!=\!14), ID-3 (n=18n\!=\!18), and ID-4 (n=0n\!=\!0). Our qualitative analysis suggests that even when a comment is valid, the PR may still be abandoned because the overall approach “doesn’t solve the root cause”. Therefore, we focused on merged PRs in our main results to ensure we capture actual code improvements that successfully transitioned into the project’s codebase.
V-B2 Analysis of Feedback on Unaddressed AI Comments
To capture impact beyond immediate code changes, we analyzed developer replies to AI comments that were not labeled as “Valid, Addressed” in Table IX. We identified a total of 50 replies and manually categorized them into four types:
  • Defending the Change (64%): Developers defended their current implementation as correct or preferred and explained why the AI suggestion was not adopted, e.g., “I don’t want to bring back ‘GetParameter’ as using type parameters is a lot nicer” (link).
  • Prompting Reflection (14%): AI comments triggered technical debates, which may lead to code changes depending on the outcome. For instance, a developer questioned the suggested optimization: “Is ‘iterrows’ really more efficient than vectorized operations? The golden rule is benchmarks” (link).
  • Assessing Tool Capabilities (12%): Beyond replying to a specific suggestion, developers evaluated the overall performance of the AI or action. For instance, one developer noted it was “interesting to learn where AI stands” (link) despite the specific suggestion being of limited value.
  • Willingness to Follow (10%): Developers expressed an intention to apply the suggestion in future iterations, e.g., “Templates could be enhanced later on to use markdown instead of raw text” (link).
Specifically, Willingness to Follow (10%) and Prompting Reflection (14%) imply that even if not addressed immediately, a comment may still trigger code changes in the future. These findings motivate future work on impact beyond code changes (e.g., guiding developer reasoning or shaping long-term code quality) to better understand the contribution of AI code reviewers to software engineering practice.
V-C Recommended Features for Effective AI Code Review
@@
  • Provide richer context to the LLM: Supplying the PR title, description, summaries of changes, or even user’s feedback helps generate more useful review comments.
@@
V-D Implications for AI Code Review Adopters
@@
Clarify the role of AI code review: Even the best-performing AI tool achieved only a 19.2% addressing rate, far below the human reviewers’ 60%. Therefore, for now, AI code review is better framed as a complement to human review, not a replacement. This requires understanding AI code review’s capability boundaries and deploying it in appropriate scenarios. Based on our LDA topic analysis (RQ3), AI reviewers are good at providing defensive programming suggestions, and developers are relatively more responsive to their comments on error handling. Accordingly, we recommend positioning AI reviewers as defenders: tightly scope them to this mandate (with targeted prompts and other recommended features), thereby suppressing category-sprawling yet low-quality comments. This allows human reviewers to focus on higher-level issues, such as overall architecture (e.g. “Try sorting at the DB level, it will be 100x faster”), development context (e.g., “Remove this as project requires Python $\geq$ 3.11”), and business logic (e.g., “This will cause issues on ARM64”). Mapping the boundaries of AI code review’s effectiveness is also an ongoing task for researchers.
@@ VI Threats to Validity
Internal Validity: In RQ2, we extracted 150 comments from over 22,000 AI-generated and 1,166 human-authored comments and obtained gold-standard labels via multi-rater annotation. To assess representativeness, the first author examined an additional 250 samples and observed similar addressing patterns as reported in our main findings. Therefore, we believe the gold standard reflects the broader dataset. For RQ3, our addressing analysis focuses on four popular actions, which may limit the generalizability of our findings. However, within the scope of our dataset, usage was highly concentrated on these four actions, making them representative of real-world deployments observed in our study. Moreover, the final analysis uses a binary classification of whether comments were addressed, which does not capture the degree or impact of the resulting code changes. Nevertheless, our evaluation shows substantial inter-rater agreement and strong LLM performance on this task, ensuring the reliability of our findings. Our feature engineering for impact factor analysis may also be incomplete or miss latent feature interactions. However, our features was comprehensively designed across four dimensions and inspired by prior works [39, 12]. Another threat is that different interpretation methods may yield different analysis results. Since our feature set spans four dimensions, we chose Random Forest to better model the complex and non-linear interactions between features, rather than the simpler model like logistic regression, which assumes a linear relationship between features. We also tried logistic regression in our experiments and it achieved a lower performance (86.1% accuracy and 0.826 Macro-F1 in 5-fold cross-validation) than our Random Forest model (88.6% and 0.846 respectively). This model choice also aligns with work [39, 12]. And the use of SHAP is a standard approach for interpreting such models as well. Therefore, our interpretation pipeline is well-justified. We also acknowledge that these interpretations describe associations between features and the predicted probability of ”Addressed”, not causal effects. If causal identification is desired, further studies (e.g., controlled experiments or developer interviews) would be needed.
External Validity: Our study includes 16 AI-based code review actions identified through a systematic Marketplace search and manual verification. Rather than sampling a subset, we included all actions that met our relevance criteria and collected the complete set of observable review comments generated by these tools in mature repositories. Although four tools account for the majority of comments (most notably anc95/ChatGPT-CodeReview), this concentration reflects actual ecosystem usage patterns at the time of data collection. We acknowledge that less popular or newly released tools are underrepresented in absolute volume. However, their limited usage inherently results in fewer observable review behaviors. Thus, our dataset captures the near-complete population of AI-based review activity occurring on GitHub Actions at the time of the study, particularly among tools with meaningful adoption. To further verify dataset coverage, we conducted a follow-up Marketplace search on July 23, 2025. We observed no substantial changes among the top-ranked actions. The only notable newcomer, hustcer/deepseek-review (ranked #7, 324 stars), had its first commit on January 29, 2025, i.e., after our action selection date (January 7, 2025). This confirms that our dataset accurately reflects the state of the ecosystem during the study period. Consequently, our findings, such as the benefits of hunk-level granularity and concrete code suggestions, are derived from analysis across multiple widely used tools and human reviews, rather than from a single-tool artifact. We will continue to monitor the evolution of the AI-based code review ecosystem in future work. Second, to mitigate the impact of inactive repositories which may not yet have established stable review workflows, we followed prior research [41] and collected data only from repositories with at least 50 PRs. However, this threshold does not fully capture the diversity of project scales. Since our dataset primarily consists of small- to medium-sized projects (typically with $\leq$ 50 non-bot contributors), our findings may not generalize to very large-scale industrial systems. Moreover, our main analysis focuses on review comments generated by file-level and hunk-level review actions and examines whether they are addressed in merged PRs. We did not consider PR-level actions. This design choice simplifies the analysis by ensuring that we can reliably observe whether comments have been addressed, but it may also narrow the generalizability of our findings. To broaden our understanding beyond this core scope, we additionally examined a sample of PR-level review comments and conducted the supplementary analyses on closed PRs and feedback on unaddressed AI comments, presented in Section V-B. Additionally, our analysis focused on English-language comments. As described in RQ2 Phase I, we began with 16,762 comments from 100 repositories. A language filter reduced this to 4,229 comments from 53 repositories, excluding about 75% of comments and 47 repositories. Among them, 32 repositories were excluded because they only contained Korean comments (e.g., 1,828 comments from KEA-DoKebi/dalkom-backend). Despite this reduction, we retained over half of the original projects. And given that English serves as the main communication language in global development activities [50], with even non-native speakers commonly adopting it, we believe our findings remain relevant to the broader open-source community.
@@ VII Related Work
Recent work explored how LLMs can be more effectively applied to code review, e.g. by incorporating pull request context [53] or through mixture-of-prompts architectures [37]. Meanwhile, benchmark-oriented studies also developed more realistic evaluation settings for AI-based code review [52, 38]. Additionally, Cihan et al. studied the impact of AI-assisted code review in industrial settings [7].
Although some studies include automatic and human evaluations of generated comments [46], most of this research has focused on comment quality rather than downstream impact. Our study complements these efforts by analyzing how AI-generated comments are used in open-source development workflows, whether they lead to code changes, and which factors influence their effectiveness.
@@ VIII Conclusion and Future Work
This paper presents a large-scale empirical analysis of AI-based code review actions on GitHub. We examine how these tools are adopted, whether their comments lead to code changes, and what factors influence their effectiveness. Our findings reveal that adoption is highly concentrated on four popular actions. Many AI-generated comments are not addressed, especially when they are vague or lack context. However, comments that are concise, specific are more likely to be addressed—particularly by less experienced contributors. Moreover, even when AI comments do not lead to immediate code changes, we find that they can still prompt developer reflection and inform future improvements. Based on our findings, we offer implications for code review tool designers, adopters, and GitHub.
We recommend treating AI code reviewers as collaborators with human reviewers. Future work should further enhance review quality by exploring richer context, better prompts, and multi-stage strategies to generate more specific and actionable feedback. It is also essential to clarify AI’s capabilities to identify optimal ways for human-AI collaboration. We hope our findings contribute to the development of more practical and developer-friendly AI-based code review tools.
Acknowledgments
This work is supported by the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM118), the National Natural Science Foundation of China (No. 62572237, No. 62302210, No. 72371125), the 2024 Development and Testing Tools Project (CEIEC-2024-ZM02-0066), the Frontier Technologies R&D Program of Jiangsu (No. BF2024059), and the Natural Science Foundation of Jiangsu Province (No. BK20241195).